OpenStudy (anonymous):

im I correct? are these all considered polynomial? 2x^2-15x+10 8x-2^-2 9x^-3+8

Parth (parthkohli):

HINT: Do they have a degree equal to or more than \(1\)?

Parth (parthkohli):

And also, positive integers?

Parth (parthkohli):

Or I mean, the lowest power greater than or equal to 0.

OpenStudy (anonymous):

yes they all do ?

terenzreignz (terenzreignz):

So much words... polynomials can't have unknowns in the numerator with negative exponents...

terenzreignz (terenzreignz):

So, butterfly... some of these aren't quite polynomials ;) Have you pinpointed them?

OpenStudy (anonymous):

terenz im confused lol i think this is the only one that is 2x^2-15x+10

Parth (parthkohli):

(Y)

terenzreignz (terenzreignz):

Well, stay your confusion, because you're RIGHT :D The second one has a negative exponent So does the third.

Parth (parthkohli):

:')

terenzreignz (terenzreignz):

NOTE: Just because all exponents are positive doesn't guarantee polynomial-ness ;) Case in point... \[\huge x^\frac52+x^2+x+1\]Is not a polynomial. Polynomials must have positive integers as exponents (in the numerator, of course)

OpenStudy (anonymous):

and these are not polynomials because they have division correct ? 2m^4-r^3/5 5x^2y^3/z

terenzreignz (terenzreignz):

@ParthKohli Constants are polynomials. So degree may be zero ;)

Parth (parthkohli):

Yeah, see the third post. The first and second were wrong :-)

terenzreignz (terenzreignz):

@ParthKohli TL;DR LOL Just kidding, I only saw that now :D Sorry, my bad @mariposa007 Well, when division is involved, it MAY still be a polynomial, provided there are no unknowns (variables) in the denominator that are raised to a positive exponent.

Parth (parthkohli):

Here's how think about it.\[\large x^{\frac{5}{2}} = x^{5 \times 2^{-1}}\]Negative power! But again.\[\large x^2 = x^{-2 \times -1}\]

Parth (parthkohli):

OK, just follow Terence. I don't wanna confuse you =_=

terenzreignz (terenzreignz):

@ParthKohli And I thought I had the tendency to complicate things.. Hey, Miss Butterfly... @mariposa007 you have an idea now?

Parth (parthkohli):

LOL

Parth (parthkohli):

@terenzreignz I love you, man.

Parth (parthkohli):

Nooooooo @ajprincess. Not me!

terenzreignz (terenzreignz):

I know. I love me too :) LOL You know, @ParthKohli I really don't know how to respond to that, so take this alpha... \[\huge \alpha\]

Parth (parthkohli):

\[\Huge \alpha\]I have a bigger alpha for you. :-)

terenzreignz (terenzreignz):

I do love alphas...

OpenStudy (anonymous):

guys guys focus LOL you have lost me. i know you might think is simple but if i am correct then 2m^4-r^3/5

Parth (parthkohli):

That's a polynomial.

OpenStudy (anonymous):

is a poly/ correct?

Parth (parthkohli):

Yupppp.

terenzreignz (terenzreignz):

Yes. Precisely, that is still a polynomial, because it's basically \[\huge \frac25m^4-\frac15r^3\]

terenzreignz (terenzreignz):

Nice reasoning, Miss Butterfly :D

Parth (parthkohli):

It belongs to the polynomial ring \(\mathbb{Z}[m]\).

Parth (parthkohli):

:3

OpenStudy (anonymous):

but not this one . this is not a poly 5x^2y^3/z

Parth (parthkohli):

Correct-o!

terenzreignz (terenzreignz):

@ParthKohli What are you doing? @ParthKohli STAHP @mariposa007 Yeah, because that is basically \[\huge 5x^2y^3z^{-1}\]

Parth (parthkohli):

;-P

terenzreignz (terenzreignz):

And you have a negative exponent...

Parth (parthkohli):

Hehe, I was having fun. :-)

OpenStudy (anonymous):

you guys have been awesome!!!

terenzreignz (terenzreignz):

Yeah. You too, Miss Butterfly :P The name's Terence ;)

OpenStudy (anonymous):

Thanks Terence!!!

Parth (parthkohli):

Say thanks to the sidekick too.

Parth (parthkohli):

:-)

terenzreignz (terenzreignz):

She already did :P Have a nice day :D

OpenStudy (anonymous):

@ParthKohli Thanks LOL you both have been awesome how can i follow you guys ? i am new to this site

Parth (parthkohli):

You can become a fan of Terence.

terenzreignz (terenzreignz):

Follow me? You mean like a stalker? Dont' you dare -.- LOL just kidding :) Just enjoy this site :D @ParthKohli You're so sweet :)

OpenStudy (anonymous):

LOL !!!

Parth (parthkohli):

@terenzreignz I love you too. :-)

terenzreignz (terenzreignz):

I'm all out of alphas, take this tau, instead... \[\Huge \tau\]

Parth (parthkohli):

<3

Parth (parthkohli):

I give you this lambda.\[\Huge \lambda\]

terenzreignz (terenzreignz):

Wait a sec... what do you mean by \[\large \mathbb{Z}[m]\]@ParthKohli For one thing, m isn't the only unknown, and even if it is, the coefficients aren't integers ಠ.ಠ LOL by the way, @mariposa007 ignore this ;)

Parth (parthkohli):

OMG, wait lol, I thought the \(r\) was \(m\) -_-

Parth (parthkohli):

What if I tell you \(\large \mathbb{Z}[m,r]\)?